what is 31 3/8th's divided by 2?
Answer by Alan3354(69443)

what is 31 3/8th's divided by 2?
-------------
3/8 is a fraction
3 8ths is a fraction
3/8th's is at best redundant.
---------
If you meant 31 3/8 divided by 2:
= 31/2 + 3/16
= 15 + 1/2 + 3/16
= 15 + 8/16 + 3/16
= 15 11/16
---------------
NOT 15 11/16th